### v1

# Understanding Addition in Transformers

Abstract
Understanding the inner workings of machine learning models like Transformers is vital for their safe and ethical use. This paper presents an in-depth analysis of a one-layer Transformer model trained for integer addition. We reveal that the model divides the task into parallel, digit-specific streams and employs distinct algorithms for different digit positions. Our study also finds that the model starts calculations late but executes them rapidly. A rare use case with high loss is identified and explained. Overall the model’s algorithm is explained in detail. These findings are validated through rigorous testing and mathematical modeling, contributing to the broader works in Mechanistic Interpretability, AI safety, and alignment. Our approach opens the door for analyzing more complex tasks and multi-layer Transformer models.
1 Introduction
Understanding the underlying mechanisms of machine learning models is essential for ensuring their safety and reliability [Barez et al., 2023, Olah et al., 2020b]. Specifically, the sub-field of mechanistic interpretability within machine learning interpretability aims to dissect the behavior of individual neurons and their interconnections in neural networks [Räuker et al., 2022]. This pursuit is part of a larger endeavor to make the decision-making processes of complex machine learning models transparent and understandable. Although models like Transformers have shown remarkable performance on a myriad of tasks, their complexity makes them challenging to interpret. Their multi-layered architecture and numerous parameters make it difficult to comprehend how they derive specific outputs [Vig, 2019]. Further, while simple arithmetic tasks like integer addition may be trivial for humans, understanding how a machine learning model like a Transformer performs such an operation is far from straightforward.
In this work, we offer an in-depth analysis of a one-layer Transformer model performing integer addition. We show that the model separates the addition task into independent digit-specific streams of work, which are computed in parallel. Different algorithms are employed for predicting the first, middle, and last digits of the answer. The model’s behavior is influenced by the compact nature of the task and the specific format in which the question is presented. Despite having the opportunity to begin calculations early, the model actually starts later. The calculations are performed in a time-dense manner, enabling the model to add two 5-digit numbers to produce a 6-digit answer in just 6 layers (See Fig. 1). A rare use case with high loss was predicted by analysis and proved to exist via experimentation. Our findings shed light on understanding and interpreting transformers. These insights may also have implications for AI safety and alignment.
As evidenced in appendices 11 and 12, our results demonstrate the model’s unique approach to integer addition across various digit lengths. The theoretical framework we develop provides a mathematical justification for the model’s behavior, substantiating our empirical observations and offering a foundation for future work.
Our main contributions are:
  • Reformulation of the traditional mathematical rules of addition into a framework more applicable to Transformers.
  • Detailed explanation of the model’s (low loss) implementation of the addition algorithm, including the problem and model constraints that informed the algorithm design.
  • Identification of a rare use case where the model is not safe to use (has high loss), and explanation of the root cause.
  • Demonstration of a successful approach to elucidating a model algorithm via rigorous analysis from first principles, detailed investigation of model training and prediction behaviours, with targeted experimentation, leading to deep understanding of the model.
Below, we provide an overview of related work (§3), discuss our methodology (§4), describe our mathematical framework (§5), our analysis of model training (§6) and model predictions (§7). We conclude with a summary of our findings and directions for future research (§8).
Figure 1: Illustration of the transformer model’s attention pattern when adding two 5-digit integers. The model attends to digit pairs sequentially from left to right, resulting in a "double staircase" pattern across layers. A: The 5 digit question is revealed token by token. The “10s of thousands" digit is revealed first. B: From the “=" token, the model attention heads focus on successive pairs of digits, giving a “double staircase" attention pattern. C: The 3 heads are time-offset from each other by 1 token such that, in each layer, data from 3 tokens is available. D: To calculate A3, the 3 heads do independent simple calculations on D3, D2 and D1. The results are combined by the MLP layer using trigrams. A3 is calculated one token before it is needed. This approach applies to all answer digits, with the first and last digits using slight variations of the approach.
2 Background
We focus on a single-layer transformer model with a vocabulary of size $V$ containing the symbols “0” to “9", “+" and “=". The model converts the human readable input (e.g. “12345+67890=") into an input sequence $(\texttt{x}_{1},\ldots,\texttt{x}_{p})$ where each $\texttt{x}_{i}\in\{1,\ldots,V\}$. Tokens are mapped to $d\_e$-dimensional embeddings by selecting the $\texttt{x}_{i}$-th column of $E\in\mathbb{R}^{d_{e}\times V}$. The model processes the input tokens, using a mechanism called “self-attention". Each input token is passed through a self-attention mechanism that calculates weighted relationships between all input tokens - capturing the importance of each token relative to others. The model then aggregates these weighted representations to produce contextually enriched representations for each token. These enriched representations are subsequently fed through feedforward neural networks (i.e. an MLP) to refine their information. Finally, the output tokens are generated based on the refined representations, and converted back to human readable format using the vocabulary (e.g. “12345+67890=80235").
3 Related Work
Interpreting and reverse engineering neural networks and transformers to find meaningful circuits has been an area of active research. Olah et al. [2020a] argued that by studying the connections between neurons and their weights, we can find meaningful algorithms (aka Circuits) in a “vision” neural network.
Elhage et al. [2021] extended this approach to transformers, conceptualizing their operation in a mathematical framework that allows significant understanding of how transformer operate internally. Tools (such as explained in Foote et al. [2023], Conmy et al. [2023b]) use this framework to semi-automate some aspects of reverse engineering. Nanda et al. [2023] reverse-engineered modular addition (e.g. 5 + 7 mod 10 = 2) showing the model used discrete Fourier transforms and trigonometric identities to convert addition to rotation about a circle.
Further, Nanda and Lieberum [2022] have argued models comprise multiple circuits, evidenced by distinct training loss curves per subtask. However, the underlying circuits remained uninvestigated. This work extends such efforts by detailing identified circuits.
From a circuit analysis lens, studies like extract graphical circuit representations and analyze component interactions. To enable analysis and interpretability, techniques in works like symbolically reverse-engineer networks by recovering computational expressions.
Research including Seth [2005] advocate analyzing networks causally, introducing structural causal models to infer mechanisms. Examinations of sequence models like have analyzed the emergence and interaction of modular components during training.
Evolutionary perspectives such as Miikkulainen [2021] elucidate how selection pressures shape hierarchical representations.
Information bottleneck analyses including relate bottlenecks to abstractions and modularization arising from forced compression. Surveys like overview techniques to disentangle explanatory factors into separate latent dimensions.
Novel objectives proposed in works like Conmy et al. [2023a] improve interpretability by encouraging modularity and disentanglement.
Figure 2: Training loss curves per digit position for a 5-digit integer addition task, showing the model trains each digit semi-independently.
4 Methodology
The integer addition problem space is very dense. For 5 digit addition, there are 10 billion distinct questions (e.g. “54321+77779=") . The model must predict all 6 answer digits correctly to get the one right answer out of 200,000 possibilities. Changing a single digit in the question changes 1 to 6 digits in the answer. The full question is only revealed one token (the “=") before the model must predict the first answer digit.
Our model was trained on 1.8 million out of 10 billion questions. After training, the model predicts answers to questions with low loss, showing the model does not rely on memorisation of training data.
Fig. 2 shows the model trains each digit semi-independently suggesting the model performs integer addition by breaking down the task into parallel digit-specific streams of computation.
The traditional human addition process first sums the units before moving on to higher value digits. This is the simplest process but relies on being able to choose the order to process the digits in. Transformer models always process text from left to right. So the model sees the higher value digits (e.g. thousands) of the question before the lower value digits (e.g. units). It doesn’t use the traditional process.
A key component of addition is the need to sum each digit in the first number with the corresponding digit in the second number. Transformer models contain “attention heads" and they are the only computational sub-component of a model that can move information between positions (aka digits or tokens). Visualising which token(s) each attention head focussed on in each layer of the calculation provided insights. While our model works with 2, 3 or 4 attention heads, 3 attention heads gives the most easily interpreted attention patterns. Fig. 3 shows the attention pattern for a single 5 digit addition calculation using 3 attention heads. Appendix 12 shows the same pattern for 10 and 15 digit addition. Appendix 12 shows the pattern with 2 or 4 attention heads.
Figure 3: The attention pattern, for a model with 3 attention heads, performing a single 5 digit addition. The pattern is 18 by 18 squares (as 54321+77779=132100 is 18 tokens). Time proceeds vertically downwards, with one additional token being revealed horizontally at each layer, giving the overall triangle shape. After the question is fully revealed (at layer 11), each head starts attending to pairs of question digits from left to right (i.e. high-value digits before lower-value digits) giving the “double staircase" shape. The three heads attend to a given digit pair in three different layers, giving a time ordering of heads.
While it’s clear the model is calculating answer digits from highest value to lowest value, using the attention heads, it’s not clear what calculation each attention head is doing, or how the attention heads are composed together to perform addition.
5 Mathematical Framework
To help investigate, we created a mathematical framework describing what any algorithm must do if it is to perform addition correctly.
Our intuition is that the model a) incrementally discovers a necessary and sufficient set of addition sub-tasks (minimising complexity), b) discovers these sub-tasks semi-independently (maximising parallelism), and c) treats each digit semi-independently (more parallelism). Our framework reflects this.
To explain the framework, let us first define $x$ and $y$ be two $n$-digit integers that need to be added, represented as vectors where $x=(x_{0},x_{1},\ldots,x_{n-1})$ and $y=(y_{0},y_{1},\ldots,y_{n-1})$.
We assert that the framework utilizes three base functions that operate on individual digit pairs. The first is Base Add (aka BA ), which calculates the sum of two digits $x_{i}$ and $y_{i}$ modulo 10, ignoring any carry over from previous columns. The second is Make Carry 1 (aka MC1 ), which evaluates if adding digits $x_{i}$ and $y_{i}$ results in a carry over of 1 to the next column. The third is Make Sum 9 (aka MS9 ), which evaluates if $x_{i}+y_{i}=9$ exactly.
In addition, the framework uses two compound functions that chain operations across digits. The first is Use Carry 1 (aka UC1 ), which takes the previous column’s carry output and adds it to the sum of the current digit pair. The second is Use Sum 9 (aka US9 ), which propagates (aka cascades) a carry over of 1 to the next column if the current column sums to 9 and the previous column generated a carry over. US9 is the most complex task as it spans three digits. For some rare questions (e.g. 00555 + 00445 = 01000) US9 applies to up to four sequential digits, causing a chain effect, with the MC1 cascading through multiple digits. This cascade requires a time ordering of the US9 calculations from lower to higher digits.
These tasks occur in the training data with different, predictable frequencies (e.g. BA is common, US9 is rarer). Compound tasks are reliant on the base tasks and so discovered later in training. The discovery of each task reduces the model loss by a different, predictable amount (e.g. BA by 50%, US9 by 5%). Combining these facts give an expected order of task discovery during training.
Figure 4: The mathematical framework (our method) predicts that during training, tasks are learnt for each digit independently, progressively increasing per digit accuracy (i.e. decreasing loss) shown as percentages). Mathematical rules cause dependencies between digits, giving an predicted ordering for perfect (i.e. zero loss) addition. The chain of blue squares relate to questions like 99999 + 00001 = 100000 where the MC1 in digit 0 causes US9 cascades through multiple other digits.
6 Training Analysis
Fig. 2 shows the model trains each digit semi-independently. Armed with the mathematical framework, we investigated each digit separately.
The Digit 0 calculation is the least interesting as it only uses BA (not UC1 or US9 ). Once discovered, Digit 0 always quickly refines to have the lowest loss and least noise (as expected). Refer Appendix 11 for graphs.
For the other digits, we categorised the training data into 3 non-overlapping subsets aligned to the BA , UC1 and US9 tasks, and graphed various combinations, finding interesting results.
The US9 graphs are much noisier than other graphs (Fig. 5). We found that the model has low loss on simple US9 cases (e.g. 45 + 55 = 100) but has high loss on US9 cascades (e.g. 445 + 555 = 1000) where the MC1 must be propagated "right to left" two 2, 3 or 4 columns. The model can’t perform these rare use cases safely, as it has a "left to right" algorithm.
Figure 5: High variability in the per digit training loss for US9 cases caused by the model’s inability to reliably do cascading US9 cases such as 445+555=1000.
Graphing the BA and UC1 use cases side by side for any one of the Digits 1, 2 and 3 shows an interesting pattern (Fig. 6). In Phase 1, both tasks have the same (high) loss. In Phase 2, both curves drop quickly but the BA curve drops faster than the UC1 curve. This “time lag" matches our expectation that the BA task must be accurate before the UC1 task can be accurate. In Phase 3, both tasks’ loss curve decrease slowly over time.
Figure 6: Training loss for digit 3 showing that, in Phase 2, the refining of Use Carry 1 lags behind Base Add . Base Add and Use Carry 1 are refined separately and have separate calculation algorithms.
Both the BA and UC1 tasks need to move data between tokens, and so will be implemented in attention head(s). Fig. 6 shows they are trained semi-independently. We choose the number of attention heads in our model with the clearest separation of tasks in the attention pattern. We find (later) that our model has separate attention heads for the BA and UC1 tasks.
Digit 4, the highest question digit, has a significantly different loss curve (shown in Fig. 7) than Digits 1, 2 and 3. This is partially explained by Digit 4 only having simple US9 cases (i.e. no US9 cascades). This does not explain the BA or UC1 differences. This difference persists with different seed values, and with 10 or 15 digit addition. We explain this difference later.
Figure 7: Training loss for digit 4 starts and stays lower for all tasks than it does for digits 1, 2 and 3. Digit 4 has a different calculation algorithm from digits 1, 2 and 3.
7 Prediction Analysis
During model prediction, we overrode (mean ablated) the model memory (residual stream) at each layer, and confirmed that the addition algorithm does not use any data generated in layers 0 to 10 inclusive. In these layers the model has not yet seen the full question and every digit in the question is independent of every other digit, making accurate answer prediction infeasible. The model also does not use the last (17th) layer. Therefore, the addition is started and completed in 6 layers (11 to 16). Further (ablation) experiments confirmed that the A0 to A4 answers are calculated one layer before being revealed.
The model has slightly different algorithms for the first digit pairs, the middle digit pairs and the last digit pairs. Fig. 1 has a simplified version of how the model calculates the middle digit pair A3. Fig. 8 has more details. For 5 digit addition, there are 2 middle digit pairs (A3 and A2) whereas for 15 digit addition there are 12 middle digit pairs.
Figure 8: A: For A3, the addition algorithm must combine information from digits 3, 2 and 1. B: The 1st head calculates MC1 on digit 1. C: The 2nd head calculates MC1 and MS9 (at most one of which can be true) on digit 2. D: The 3rd head calculates Base Add on digit 3. E: The MLP layer uses trigrams to combine the information from the 3 heads to give the final answer A3, one layer (token) before it is output.
The A3 addition algorithm has three clauses related to digits 3, 2 and 1. Ablating each head in turn shows that the 3rd head has most impact on loss, the 2nd head has less impact, and the 1st head has little impact. This aligns with the intuition that the sum “D3 + D3’" matters most, the MC1 from the previous digit (D2 + D2’) matters less, and the rare MC1 from the previous previous digit (D1 + D1’) matters least.
The last two digits, A1 and A0, use a simplified a version of the A3 algorithm, with some of the three clauses not needed.
The A3 algorithm could also successfully be applied to A4. But the Digit 4 training curve is better (faster) than the middle digits. The attention patterns shows that for A4, the model is using all the heads in layer 11 (the “=" token) when the A3 algorithm doesn’t require this. Uniquely, A4 utilises more “compute" than is available to A3, A2, A1 or A0. We assume the model uses this advantage to implement a faster-training and lower-loss algorithm for A5 and A4. We haven’t worked out the details of this.
Mean ablating the 1st or 2nd head slightly increased the average loss for BA questions from 0.05 to 0.08, whereas ablating the 3rd head substantially increased the loss to 3.7, confirming that the 3rd head is doing the BA task.
The MLP can be thought of as a “key-value pair" memory [Meng et al., 2022, Geva et al., 2021] that can hold many bigrams and trigrams. We claim our MLP pulls together the two-state 1st head result, the tri-state 2nd head result and the ten-state 3rd head result value, treating them as a trigram with 60 (2 x 3 x 10) possible keys. For each digit, the MLP has memorised the mapping of these 60 keys to the 60 correct digit answers (0 to 9). We haven’t proven this experimentally. Our MLP is sufficiently large to store this many mappings with zero interference between mappings [Elhage et al., 2022].
Despite being feasible, the model does not calculate the task MC1 in layers 7 to 11. Instead it completes each digit calculation in 1 layer, possibly because there are training optimisation benefits in generating a “compact" algorithm.
This algorithm explains all the observed prediction behaviour - including the fact that the model can calculate a simple US9 case but not a cascading US9 case. We assume that, given the dense nature of the question and answer, and the small model size, the model does not have sufficient time and compute resources to implement both UC1 and US9 accurately, and so preferences implementing the more common (UC1 ) case, and only partially implements the more complex and rare (US9 ) case.
8 Conclusions and Future Work
This work demonstrates a successful approach to reverse engineering and elucidating the emergent algorithm within a transformer model trained on integer addition. By combining mathematical analysis, empirical investigation of training and prediction, and targeted experimentation, we are able to explain how the model divides the task into parallel digit-specific streams, employs distinct subroutines for different digit positions, postpones calculations until the last possible moment yet executes them rapidly, and struggles with a specific rare case.
Our theoretical framework of necessary addition subtasks provides a foundation for the model’s behavior. The digit-wise training loss curves reveal independent refinement consistent with separate digit-specific circuits. Attention patterns illustrate staging and time-ordering of operations. Controlled ablation experiments validate ouur hypothesis about algorithmic elements’ roles. Together these methods enable a detailed accounting of the model’s addition procedure.
This methodology for mechanistic interpretability, when applied to broader tasks and larger models, can offer insights into not just what computations occur inside complex neural networks, but how and why those computations arise. Such elucidation will be increasingly important for ensuring the safety, reliability and transparency of AI systems.
Our study paves the way for numerous potential research avenues. Recognizing the challenges in rare cases can inspire methods to enhance the robustness of addition models. The established framework might be adapted to elucidate models for integer subtraction or multiplication. By integrating proven and effective addition modules into a larger, untrained network geared towards multiplication, the training process could be expedited. Further, decoding the multiplication algorithm becomes more straightforward when the addition-related tasks are already recognized and deemed reliable. Utilizing this modular approach can simplify the understanding of intricate algorithms, propelling advancements in the field of mechanistic interpretability. In summary, this research underscores that diving deep into the workings of contemporary machine learning can highlight valuable strengths, pinpoint areas for improvement, and present avenues for accelerated progress.
9 Reproducibility Statement
To ensure our work is reproducible, we provide the full source code in the supplementary materials, as well as all necessary data and parameters and instructions to reproduce our experimental results.